# Looking Backward: Retrospective Backward Synthesis for Goal-Conditioned GFlowNets

### v1

@@ 5.2 Bit Sequence Generation
In this section, we investigate the performance of RBS-GFN in the bit sequence generation task (Malkin et al., 2022). Unlike previous approaches that generate these sequences in a left-to-right manner (Malkin et al., 2022, Madan et al., 2023a), we adopt a non-autoregressive prepend/append Markov decision process following Shen et al. (2023). The action space includes pretending or appending a $k$-bit word from the vocabulary $V$ to the current state, which increases the difficulty of the task (as the underlying structure of the problem is a directed acyclic graph rather than a simple tree (Malkin et al., 2022)). We consider bit sequence generation with small, medium, and large sizes with increasing lengths and vocabulary sizes following Pan et al. (2023a).
@@
As shown in Fig. 9, even the strongest OC-GAFN method struggles to learn efficiently given a limited training budget, while all other baselines completely fail. In contrast, RBS-GFN achieves high success rates of approximately 100% with fast convergence across different scales of the tasks. It is worth noting that RBS-GFN without either age-based sampling (denoted as RBS-GFN-w/o-AgeSample) or intensified reward feedback (see detailed discussion in Appendix A.1) both fail to generalize to more complex tasks, including medium and large, demonstrating the importance of our proposed techniques in enabling RBS-GFN to efficiently learn across various levels of complexity.
@@ 6 Conclusion
In this paper, we address the critical challenges of realizing goal-directed behavior and learning in GFlowNets. To overcome the training challenge due to extremely sparse rewards, we propose a novel method called Retrospective Backward Synthesis, which significantly improves the training of goal-conditioned GFlowNets by synthesizing backward trajectories. Our extensive experiments demonstrate state-of-the-art performance in terms of both success rate and generalization ability, which outperforms strong baselines. For future work, it is promising to further improve our method, e.g., sampling method considering alternative priorities (Sujit et al., 2023).
@@ Appendix A Intensified Reward Feedback
$\mathcal{L}_{\rm GC-GFN}=\left(\log{F_{\theta}(s|y)P_{F}(s^{\prime}|s,y,\theta)}-\log{\left[CR(x,y)P_{B}(s|s^{\prime},y,\theta)\right]}\right)^{2},$ (7)
which can be degenerated to Eq. (5) when we set $C=1$. In order to enable effective reward propagation, we set C as a large value in real implementation. Below we show a large value of $C$ can scale the gradient w.r.t. $P_{B}$ without influencing $P_{F}$ and $F$.
$\frac{\partial\mathcal{L}_{\rm GC-GFN}}{\partial\theta}$ $=2\left(\log F_{\theta}(s|y)+\log P_{F}(s^{\prime}|s,y,\theta)-\log\left[C\times R(x,y)P_{B}(s|s^{\prime},y,\theta)\right]\right)\times$ (8)
$\Bigl{(}\frac{1}{F_{\theta}(s|y)}\frac{\partial F_{\theta}(s|y)}{\partial\theta}+\frac{1}{P_{F}(s^{\prime}|s,y,\theta)}\frac{\partial P_{F}(s^{\prime}|s,y,\theta)}{\partial\theta}$
$-\frac{1}{C\times R(x,y)P_{B}(s|s^{\prime},y,\theta)}\frac{\partial(C\times R(x,y)P_{B}(s|s^{\prime},y,\theta))}{\partial\theta}\Bigr{)}$
$=2\left(\log F_{\theta}(s|y)+\log P_{F}(s^{\prime}|s,y,\theta)-\log\left[C\times R(x,y)P_{B}(s|s^{\prime},y,\theta)\right]\right)\times$
$\Bigl{(}\frac{1}{F_{\theta}(s|y)}\frac{\partial F_{\theta}(s|y)}{\partial\theta}+\frac{1}{P_{F}(s^{\prime}|s,y,\theta)}\frac{\partial P_{F}(s^{\prime}|s,y,\theta)}{\partial\theta}$
$-\frac{1}{P_{B}(s|s^{\prime},y,\theta)}\frac{\partial(P_{B}(s|s^{\prime},y,\theta))}{\partial\theta}\Bigr{)}$
@@ Appendix B Hierarchical Goal Decomposition
It is still challenging to tackle problems with extremely large-scale state spaces, and even our method can fail in these scenarios. To address this problem, we propose a hierarchical method to decompose the task into several low-level tasks that are easier to complete. Leveraging the consistent sequential structure in compositional GFlowNets tasks, we can set sub-goals manually, eliminating the need to learn a sub-goal generation policy additionally (Chane-Sane et al., 2021, Kim et al., 2021, Nasiriany et al., 2019). After training sub-level policies, all the generated sub-goals can be combined together to obtain the final goal.
@@ Appendix C Experimental Setup
We build our implementation for all baselines and environments upon publicly available open-source repositories. [^1] https://github.com/GFNOrg/gflownet
@@
The GridWorld (Bengio et al., 2021) is conceptualized as a 2-dimensional hypercube with side length $H:\{(s^{1},s^{2})|s^{i}\in\{0,1,\cdots,H-1\}$, where the model learns to achieve any given goals (outcomes) starting from a fixed initial state $(0,0)$. We examine grids with $H$ set to $32$, $64$, and $128$, respectively, resulting in different levels of difficulty categorized as small, medium, and large. The agent receives a positive reward of $1$ only if it reaches the desired goal state. We use the Adam (Kingma and Ba, 2014) optimizer with a learning rate of $1e^{-3}$ for $2e^{4}$ training steps.
@@
This task requires the model to generate sequences of length $n$ by pretending or appending a $k$-bit word to the current state. We consider $k=2,n=40$ for small task, $k=3,n=60$ for medium task, and $k=5,n=100$ for large one. We use the Adam (Kingma and Ba, 2014) optimizer with a learning rate of $5e^{-4}$ for $1e^{5}$ training steps.
@@
Similar to the bit sequence generation task, the agent prepends or appends a symbol from the vocabulary with a size of $4$ to the current state at each step to generate a sequence of length $8$. We use the Adam (Kingma and Ba, 2014) optimizer with a learning rate of $5e^{-4}$ for $5e^{3}$ training steps.
@@
This biological task requires the agent to generate antimicrobial peptides (AMP) with lengths of $50$ (Jain et al., 2022) from a vocabulary with size of $20$. For both RBS-GFN, Hier-RBS-GFN and all the baselines, we use the Adam (Kingma and Ba, 2014) optimizer with a learning rate of $5e^{-4}$ for $1e^{5}$ training steps.
@@ C.2 Baselines
  • Following the implementation in (Andrychowicz et al., 2017), DQN w/ HER leverages both deep Q-learning algorithm (Mnih et al., 2013, Jang et al., 2019) and HER technique (Andrychowicz et al., 2017) to learn a near-optimal policy.
@@
Figure 12: Performance of GFN with SubTB in TF Bind tasks.
@@ Appendix D Additional Experimental Results
SAC performance in TFBind sequence generation task.
We investigate the performance of SAC, known as soft actor-critic algorithm (Haarnoja et al., 2018), which is an entropy-regularized RL method rather than standard $\arg\max$ DQN. We evaluate the performance of SAC in the TF Bind
Figure 14: Performance of SAC in TF Bind tasks.
sequence generation task. As the action space is this task is discrete, we implement a discrete SAC algorithm based on the codes from CleanRL (Huang et al., 2022). From the results shown in Fig 14, we observe that SAC (w/ HER) even performs worse than the DQN algorithm. We hypothesize that it is because SAC prefers new states to maximize the entropy rather than high-reward states to complete the task. Although HER can provide abundant of successful experiences, it is still not enough for SAC to succeed.

### v2

@@ 5.2 Bit Sequence Generation
In this section, we investigate the performance of RBS-GFN in the bit sequence generation task (Malkin et al., 2022). Unlike previous approaches that generate these sequences in a left-to-right manner (Malkin et al., 2022; Madan et al., 2023a), we adopt a non-autoregressive prepend/append Markov decision process following Shen et al. (2023), which is a more challenging task compared to the one studied in Pan et al. (2023a). The action space includes pretending or appending a $k$-bit word from the vocabulary $V$ to the current state, which increases the difficulty of the task (as the underlying structure of the problem is a directed acyclic graph rather than a simple tree (Malkin et al., 2022)). We consider bit sequence generation with small, medium, and large sizes with increasing lengths and vocabulary sizes following Pan et al. (2023a).
@@
As shown in Fig. 10, even the strongest OC-GAFN method struggles to learn efficiently given a limited training budget, while all other baselines completely fail. In contrast, RBS-GFN achieves high success rates of approximately 100% with fast convergence across different scales of the tasks. It is worth noting that RBS-GFN without either age-based sampling (denoted as RBS-GFN-w/o-AgeSample) or intensified reward feedback (see detailed discussions in Appendix A.1) both fail to generalize to more complex tasks, including medium and large, demonstrating the importance of our proposed techniques in enabling RBS-GFN to efficiently learn across various levels of complexity.
5.3 TF Bind Generation
5.4 AMP Generation
@@
5.5 Application: Downstream Finetuning
A notable advantage of GC-GFlowNets is that the pre-trained policy can be leveraged to handle downstream tasks with unseen rewards, unlike the typical fine-tuning process of reinforcement learning as they generally learn reward-maximization policies that may discard valuable information (Pan et al., 2023a). In this section, we study the application of GC-GFlowNets and validate its effectiveness for adapting to downstream bit sequence generation tasks with unseen rewards in different scales. We generally follow the experimental design in Pan et al. (2023a), while we train GC-GFlowNets with fewer pre-training budgets, which better illustrates the efficiency of our method in the pre-training stage. From the results shown in Fig. 12, we find that RBS-GFN outperforms OC-GAFN by a large margin, as a more efficient and effective pre-trained goal-reaching strategy can significantly contribute to the fine-tuning process, while OC-GAFN struggles to efficiently discover modes given limited pre-training budgets in this challenging goal-conditioned training stage.
Figure 12: Fine-tuning GC-GFlowNets on the downstream sequence generation task with different scales. We report the number of modes discovered during training. The number of modes is calculated using a sphere-exclusion procedure. A candidate is added to the list of modes if it is above a certain reward threshold and is further away than some distance threshold from all other modes.
@@ 6 Conclusion
In this paper, we address the critical challenges of realizing goal-directed behavior and learning in GFlowNets. To overcome the training challenge due to extremely sparse rewards, we propose a novel method called Retrospective Backward Synthesis, which significantly improves the training of goal-conditioned GFlowNets by synthesizing backward trajectories. Our extensive experiments demonstrate state-of-the-art performance in terms of both success rate and generalization ability, which outperforms strong baselines. For future work, it is promising to further improve our method, e.g., sampling method considering alternative priorities (Sujit et al., 2023). More discussions about our work are provided in Appendix E.
Acknowledgments
This work of Haoran He and Ling Pan is supported by National Natural Science Foundation of China 62406266. Huazhe Xu is supported by Tsinghua dushi program. We also thank the anonymous reviewers for their valuable suggestions.
Ethics Statement
This paper presents work whose goal is to advance the field of Machine Learning. Specifically, we propose a novel method called RBS to enhance the learning of GC-GFlowNets. Since this method is easy to reproduce (as we will release our code soon) and exhibits the SOTA performance, it encourages future research to further advance this field. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Reproducibility Statement
All details of our experiments can be found in Appendix C, which includes descriptions of the tasks, experimental setup, network architecture, and hyperparameters. The proof of intensified reward feedback is referred to in Appendix A. The code will be open-sourced upon publication of this work.
@@ Appendix A Intensified Reward Feedback
$\mathcal{L}_{\rm GC-GFN}=\left(\log{F_{\theta}(s|y)P_{F}(s^{\prime}|s,y,\theta)}-\log{\left[{\color[rgb]{1,0,0}C}R(x,y)P_{B}(s|s^{\prime},y,\theta)\right]}\right)^{2},$ (7)
which can be degenerated to Eq. (5) when we set ${\color[rgb]{1,0,0}C}=1$. In practice, we set ${\color[rgb]{1,0,0}C}$ to a large value to facilitate effective reward propagation. Below, we demonstrate that a large ${\color[rgb]{1,0,0}C}$ scales the gradient with respect to $P_{B}$ without affecting $P_{F}$ or $F$. We show that
$\frac{\partial\mathcal{L}_{\rm GC-GFN}}{\partial\theta}=2\times Z\frac{\partial\log Z}{\partial\theta},$ (8)
where
$Z=\log F_{\theta}(s|y)+\log P_{F}(s^{\prime}|s,y,\theta)-\log\left[{\color[rgb]{1,0,0}C}R(x,y)P_{B}(s|s^{\prime},y,\theta)\right],$ (9)
$\frac{\partial\log Z}{\partial\theta}=\frac{1}{F_{\theta}(s|y)}\frac{\partial F_{\theta}(s|y)}{\partial\theta}+\frac{1}{P_{F}(s^{\prime}|s,y,\theta)}\frac{\partial P_{F}(s^{\prime}|s,y,\theta)}{\partial\theta}-\frac{1}{{\color[rgb]{1,0,0}C}R(x,y)P_{B}(s|s^{\prime},y,\theta)}\frac{\partial\left({\color[rgb]{1,0,0}C}R(x,y)P_{B}(s|s^{\prime},y,\theta)\right)}{\partial\theta}.$
Since the scaling coefficient ${\color[rgb]{1,0,0}C}$ does not depend on the model parameters $\theta$, the derivative w.r.t. $\theta$ simplifies as follows:
$\frac{\partial\left({\color[rgb]{1,0,0}C}R(x,y)P_{B}(s|s^{\prime},y,\theta)\right)}{\partial\theta}={\color[rgb]{1,0,0}C}R(x,y)\frac{\partial P_{B}(s|s^{\prime},y,\theta)}{\partial\theta}.$ (11)
Substituting this into Eq. 10, we obtain:
[…]
where ${\color[rgb]{1,0,0}C}$ is eliminated. Consequently, ${\color[rgb]{1,0,0}C}$ only appears in the last term of $Z$. Therefore it only affects $P_{B}$, leaving gradients w.r.t. $P_{F}$ and $F_{\theta}$ unchanged.
@@ Appendix B Hierarchical Goal Decomposition
It is still challenging to tackle problems with extremely large-scale state spaces, and even our method can fail in these scenarios. To address this problem, we propose a hierarchical method to decompose the task into several low-level tasks that are easier to complete. Leveraging the consistent sequential structure in compositional GFlowNets tasks, we can set sub-goals manually, eliminating the need to learn a sub-goal generation policy additionally (Chane-Sane et al., 2021; Kim et al., 2021; Nasiriany et al., 2019). After training sub-level policies, all the generated sub-goals can be combined together to obtain the final goal.
@@ Appendix C Experimental Setup
We build our implementation for all baselines and environments upon publicly available open-source repositories. [^1] https://github.com/GFNOrg/gflownet The code will be open-sourced upon publication of the work.
@@
The GridWorld (Bengio et al., 2021) is conceptualized as a 2-dimensional hypercube with side length $H:\{(s^{1},s^{2})|s^{i}\in\{0,1,\cdots,H-1\}$, where the model learns to achieve any given goals (outcomes) starting from a fixed initial state $(0,0)$. We examine grids with $H$ set to $32$, $64$, and $128$, respectively, resulting in different levels of difficulty categorized as small, medium, and large. The agent receives a positive reward of $1$ only if it reaches the desired goal state. We use the Adam (Kingma & Ba, 2014) optimizer with a learning rate of $1e^{-3}$ for $2e^{4}$ training steps.
@@
This task requires the model to generate sequences of length $n$ by pretending or appending a $k$-bit word to the current state. We consider $k=2,n=40$ for small task, $k=3,n=60$ for medium task, and $k=5,n=100$ for large one. We use the Adam (Kingma & Ba, 2014) optimizer with a learning rate of $5e^{-4}$ for $1e^{5}$ training steps.
@@
Similar to the bit sequence generation task, the agent prepends or appends a symbol from the vocabulary with a size of $4$ to the current state at each step to generate a sequence of length $8$. We use the Adam (Kingma & Ba, 2014) optimizer with a learning rate of $5e^{-4}$ for $5e^{3}$ training steps.
@@
This biological task requires the agent to generate antimicrobial peptides (AMP) with lengths of $50$ (Jain et al., 2022) from a vocabulary with size of $20$. For both RBS-GFN, Hier-RBS-GFN and all the baselines, we use the Adam (Kingma & Ba, 2014) optimizer with a learning rate of $5e^{-4}$ for $1e^{5}$ training steps.
@@ C.2 Baselines
  • Following the implementation in (Andrychowicz et al., 2017), DQN w/ HER leverages both deep Q-learning algorithm (Mnih et al., 2013; Jang et al., 2019) and HER technique (Andrychowicz et al., 2017) to learn a near-optimal policy.
@@
C.3 Details of Offline Experiments
For offline dataset collection, we store the samples recorded in the replay buffer of vanilla GFN during training until convergence. This dataset includes 8,016 trajectories with varying levels of performance. Given that the training of GFlowNets is off-policy, we reuse the learning objective in Eq. 6 without modification. The implementation details of our offline algorithm (RBS-GFN) and the baselines are illustrated below:
  • (offline-) RBS-GFN (ours): Similar to the online version of RBS-GFN, the dataset is augmented by the reverse trajectories collected by $P_{B}$ at each training step.
  • (offline-) OC-GAFN: Unlike RBS-GFN, OC-GAFN fails to generate new trajectories. Instead, it only augments the dataset by relabeling the outcomes of failed trajectories with their actual terminal states.
  • (offline-) GFN: This baseline utilizes the original dataset without any additional data generation or augmentation processes.
To ensure a fair comparison, all other settings, including the network architecture and hyperparameters, are kept the same as those used in the online implementations.
(a) small
(b) medium
(c) large
Figure 14: Learning curves of offline experiments on three different scales of GridWorld. Success rates are averaged over three random seeds.
@@ Appendix D Additional Experimental Results
Figure 15: Additional experimental results. (a) Success rates on GridWorld with a horizon of $32$. The x-axis corresponds to the wall clock time. (b) Ablation study with different values of the decay hyperparameter $\beta$. RBS-GFN is robust to different $\beta$. (c) Success rates on RNA generation (Pan et al., 2023a). RBS-GFN consistently outperforms the strongest baseline, OC-GAFN, on the task.
D.1 Computation Overhead
RBS-GFN is efficient since we only need to synthesize a single $\tau^{\prime}$ using backward policy $P_{B}$ from the goal, and rollout a minibatch of data at each training step. It is also worth noting that the strongest baseline OC-GAFN requires training two GFN models (i.e., an unconditioned GAFN model and a goal-conditioned GFN model), while RBS-GFN only needs to train a single goal-conditioned GFN agent, which largely reduces training compute requirements. We quantify the wall-clock time and corresponding achieved success rates on the GridWorld task with a horizon of $32$. From the results shown in Fig. 15(a), we find that RBS-GFN achieves a significantly higher success rate in less time compared to the previous strongest method, OC-GAFN.
D.2 Results on the RNA Generation task
We further evaluate the performance of RBS-GFN on the RNA generation task (Lorenz et al., 2011), which involves constructing RNA sequences following Pan et al. (2023a). We compare our method with the strongest baseline, OC-GAFN. From the results shown in Fig. 15(c), we observe that RBS-GFN consistently achieves the best performance.
D.3 Robustness to Hyperparameters
Regarding the reward intensification technique, the scaling coefficient $C$ is the only task-dependent hyperparameter that requires tuning, as it depends on the nature of specific tasks and also accommodates different horizons. However, this can be easily done through standard techniques like grid search (similar to tuning conventional hyperparameters such as the learning rate (Malkin et al., 2022; Jain et al., 2022)). As for backward policy regularization, we demonstrate that the RBS-GFN exhibits robust performance across a wide range of values for the decay hyperparameter $\beta$, consistently achieving 100% success rate with high sample efficiency, as shown in Fig. 15(b).
D.4 Comparison with Model-Based Goal-Conditioned RL
To further demonstrate the effectiveness of our proposed RBS-GFN, we carefully design a sophisticated model-based GC-RL method following MHER (Yang et al., 2021) and Dreamerv3 (Hafner et al., 2023) for additional comparison. Specifically, we consider actor-critic learning introduced in Deamerv3, and employ the model-based imagination technique introduced in MHER to augment the training data. It is noteworthy that while previous model-based methods like MHER augment datasets by forward imagination based on HER, RBS introduces a novel way to sample backward trajectories to increase both data quality and diversity. The results shown in Fig. 16(a) demonstrate that RBS-GFN consistently outperforms this GC-RL method by a large margin.
D.5 Performance on Stochastic Environments
To further demonstrate that RBS-GFN can generalize to stochastic dynamics, we build it upon stochastic GFN (Pan et al., 2023d) and consider randomness in the environment following Machado et al. (2018). Specifically, the environment transitions according to the selected action with probability $1-\alpha$, while with probability $\alpha$ the environment executes a randomly chosen action. Here we take the grid environments for evaluation and set $\alpha=0.01$ to make the environments stochastic. The experimental results shown in Fig. 16(d) demonstrate that RBS-GFN can also generalize to stochastic environments and achieve higher success rates compared with the strongest baseline OC-GAFN. Given the inherent randomness in the environments, which can significantly influence goal-reaching strategies, it is reasonable for the overall performance to decline compared to that in deterministic environments.
D.6 Robustness to Different Reward Structures
To demonstrate that RBS-GFN is robust to different reward structures, we add additional experiments on the GridWorld tasks, where the agent receives dense rewards (which corresponds to an easier setting compared to the sparse reward case we studied in the main paper). Concretely, the reward is defined as the Manhattan distance between the current state and the desired goal. The results shown in Fig 16(b) demonstrate that RBS-GFN achieves even further performance improvements in the dense rewards setting.
Figure 16: Additional experimental results on the GridWorld benchmark. (a) Success rates compared with an advanced model-based goal-conditioned RL method. (b) RBS-GFN can generalize to the dense rewards structure and gain further performance improvement. (c) RBS-GFN, with the age-based sampling technique, outperforms its variant with PER. (d) Success rates on the stochastic GridWorld environments (Pan et al., 2023d) compared with OC-GAFN (across 3 random seeds). RBS-GFN consistently outperforms the strongest baseline.
D.7 Comparison with Prioritized Experience Replay (PER)
Specifically, we follow the standard PER setting (Schaul et al., 2016) with $\alpha=0.7$ and $\beta=0.4$, and we adopt the GC-GFlowNet loss instead of TD error as a priority to suit our scenario. We utilize the open-sourced codes in https://github.com/Howuhh/prioritized_experience_replay to implement it. From the results in Fig 16(c), we observe that our age-based sampling outperforms PER by a large margin that learns more efficiently. We hypothesize that this is because the GC-GFlowNet loss is not stable; for some data samples, the loss would remain irreducible and stay high throughout the training process. Consequently, PER restricts training data coverage, leading to reduced overall performance. In contrast, our age-based sampling technique ensures that experiences are leveraged more thoroughly.
D.8 Generalization and Versatility
@@
Figure 17: (a) Performance of SAC in TF Bind tasks. (b) Performance of GFN with SubTB in TF Bind tasks.
D.9 Comparison with SAC
We investigate the performance of SAC, known as soft actor-critic algorithm (Haarnoja et al., 2018), which is an entropy-regularized RL method rather than standard $\arg\max$ DQN. We evaluate the performance of SAC in the TF Bind sequence generation task. As the action space is this task is discrete, we implement a discrete SAC algorithm based on the codes from CleanRL (Huang et al., 2022). From the results shown in Fig 17(a), we observe that SAC (w/ HER) even performs worse than the DQN algorithm. We hypothesize that it is because SAC prefers new states to maximize the entropy rather than high-reward states to complete the task. Although HER can provide abundant of successful experiences, it is still not enough for SAC to succeed.
@@ D.9 Comparison with SAC
Appendix E Limitations and Discussions
E.1 Limitations and Future Work
In this paper, we mainly address key training challenges in GC-GFlowNets problems, and study standard evaluation benchmarks from the GFlowNets literature (Bengio et al., 2021; 2023) with structured tasks (e.g., DNA/RNA generation) where the dynamics are known and well-defined. For some tasks where environment dynamics might be unknown or infeasible to directly model, we can learn a backward dynamic model $f$ to predict the previous state, e.g., $s_{t-1}=f(s_{t},a_{t-1})$, following Höftmann et al. (2023); Pan et al. (2023d). With a sufficiently collected dataset, learning a dynamic model is feasible as it can be framed as a regression problem. We hope our work can inspire future research in this promising direction studying unknown dynamics in the environment.
To align with the established and commonly used benchmarks and evaluation protocols in both GFlowNets (Bengio et al., 2023) and GC-GFlowNets (Pan et al., 2023a) literature, our work primarily focuses on deterministic and discrete environments, which have been well-established and studied. While recent theoretical work (Bengio et al., 2023) explore continuous GFlowNets, their practical implementations and applications remain limited, as highlighted in Jain et al. (2023), where training continuous GFlowNets poses significant challenges and scaling them to realistic tasks remains an open problem in the field. We leave the extension of our method to continuous environments for future work.
E.2 Backward Learning in Reinforcement Learning (RL)
Previous works (Goyal et al., 2019; Edwards et al., 2018; Lai et al., 2020; Wang et al., 2021) in model-based RL leverage backward world models to optimize policies for returning to high-value states, while they fail to address the reward sparsity challenges in goal-conditioned RL. Höftmann et al. (2023) use a backward dynamic model to generate trajectories for learning goal-conditioned policies by imitation learning, which sidesteps the requirement of rewards for policy learning. However, its performance heavily relies on the quality of the learned backward model and has only been evaluated in relatively simple maze environments. Model-based goal-conditioned RL has been emerging as a promising direction, which employs a learned world model to imagine future trajectories to improve policy learning in an online (Yang et al., 2021; Charlesworth & Montana, 2020; Wang et al., 2024) or offline manner (Kim et al., 2024; Wang et al., 2023). However, these methods primarily leverage forward imagination through the learned dynamic model to improve policy learning, whereas RBS introduces a novel approach by sampling backward trajectories, thereby enhancing both data quality and diversity.